what is the sum of the measures of the interior angles of a regular 72-gon?
Answer by stanbon(75887)

what is the sum of the measures of the interior angles of a regular 72-gon?
---
sum = (n-2)180
----
sum = 70*180 = 12600 degrees
==================================
Cheers,
Stan H.